# Causal-Adapter: Taming Text-to-Image Diffusion for Faithful Counterfactual Generation

### v3

@@ Appendix A Extended Related Works
As shown in our motivational study (Appendix B), we evaluate the adaptation of text-to-image diffusion models for counterfactual image generation. Our findings show that relying solely on a frozen diffusion backbone with prompt-tuning is insufficient, as it fails to jointly represent causal semantic attributes and image embeddings. As a result, the model struggles to achieve precise counterfactual reasoning and generation. This highlights the need for an adaptive mechanism within the diffusion model, enhanced with injected causal semantics.
@@
Appendix B Full Motivational Study Results
@@
3. Can existing T2I based methods do fine-grained editing? In causal editing, interventions on a single factor (e.g., increasing object size) should consistently induce predictable and proportional changes in the generated image. However, with current T2I based methods, numerical attributes must be mapped to discrete linguistic tokens, making it difficult to express edits along a continuous range. In practice, the primary way to adjust editing strength is by tuning the classifier-free guidance scale (Ho & Salimans, 2021). As shown in Figure 12, both NTI and MCPL fail to achieve fine-grained anatomical counterfactual editing of brain ventricular volume, even with different guidance scales. Their text-only control mechanisms cannot reliably translate numeric interventions into gradual visual changes. This limitation is particularly problematic in medical imaging domains, where precise, numerically controlled edits (e.g., adjusting brain or ventricular volume in MRI scans) are essential for simulating disease progression.
@@
Appendix C Dataset and Implementation
@@ C.1 Implementation Details
Unless specified, experimental setups on CelebA and ADNI follow the benchmark protocol in (Melistas et al., 2024). To ensure fairness, we adopt the same causal mechanism (normalizing flow) as described in the benchmark implementation. Experimental setup on Pendulum follows the configuration in CausalDiffAE (Komanduri et al., 2024c). Since both approaches assume linear causal modeling, we use our constructed simple MLP-based causal mechanism.
@@
For counterfactual image generation, we employ DDIM inversion (Song et al., 2021). To best preserve identity, we set the inversion guidance scale to $1.0$ (no classifier-free guidance (CFG)). For the forward process, we use CFG: CelebA uses $\alpha=3.0$ for effective edits, while ADNI and Pendulum perform well without CFG. Optionally, we incorporate token-level attention guidance manipulation following (Ju et al., 2024), replacing only the intervened token attentions while keeping the others fixed.
@@
CelebA ADNI Pendulum
Train Set Size 162,770 10,780 5,000
Validation Set Size 19,867 0 500
Test Set Size 19,962 2,240 2,000
Resolution $\times 256\times 3 256×256$$\times 256\times 3$ $256\times 256\times 3$
Downsampled Resolution $64\times 64\times 3$ $192\times 192\times 1$ $94\times 94\times 3$
Batch Size 40 40 40
Training Steps 200k 100k 50k
Num of Attributes 4 6 4
Prompt Template
‘‘a human is @
and * and & and #’’
‘‘an MRI image is @
and * and &’’
‘‘an image is @
and * and & and #’’
CTC Temperature $\tau$ 0.2 0.2 0.2
CTC Scale $\lambda$ 0.0005 0.0005 0.0005
DDIM Steps $T$ 100 100 100
Guidance Scale $\alpha$ 3.0 1.0 1.0
@@ C.2 Metrics
Minimality evaluates whether a counterfactual differs from the factual image only in the intervened parent attribute, ideally leaving all other attributes unaffected. Counterfactual Latent Divergence (CLD) quantifies this by measuring the “distance” between counterfactual and factual images in a latent space (Sanchez & Tsaftaris, 2022). Intuitively, CLD captures a trade-off: the counterfactual should move sufficiently away from the factual class, but not farther than real samples belonging to the counterfactual class. Following Melistas et al. (2024), we compute CLD using an unconditional VAE. Specifically, we measure the KL divergence between the latent distributions of real and counterfactual images. The metric is minimized when both probabilities remain low, reflecting the balance between departing from the factual class while remaining closer to the counterfactual class than unrelated real samples.
@@ C.3 Full Regularization Algorithm
In the following, we present the fully regularized Causal-Adapter in Algorithms 34.
@@
1:Input: image $x$, semantic attributes $Y$, frozen $\{\mathcal{E},c_{\phi},\epsilon_{\theta}\}$.
2:Output: trained causal mechanisms $F$, adapter $\ddot{\epsilon}_{\psi}$ and token embeddings $V$.
3:# training causal mechanisms
4: Fit $F$ on $Y$ by minimising Eqn. 3
5:# training conditional adapter
@@
[…]
8:for $step=1$ to $T$ do
9:  Encode $z_{t}=\mathcal{E}(x,t)$
10:  $r_{t}\coloneqq\ddot{\epsilon}_{\psi}(z_{t},t,V(Y))$
11:  $\psi,\,G,\,C:=\operatorname*{arg\,min}_{{\psi,\,G,\,C}}\mathcal{L}_{\text{DM}}+\mathcal{L}_{\text{CTC}}$// CTC
12:end for
13:Return $F$, $\ddot{\epsilon}_{\psi}$ and $V$.
Algorithm 4 Causal-Adapter: Inference
1:Input: image $x$, semantic attributes $Y$, frozen $\{\mathcal{E},c_{\phi},\epsilon_{\theta}\}$, trained $\{F,\ddot{\epsilon}_{\psi},V\}$
2:Output: counterfactual image $\bar{x}$.
3:Encode $z_{0}=\mathcal{E}(x,0)$
[…]
6:for $t=0$ to T1T-1 do
7:  $r_{t}\coloneqq\ddot{\epsilon}_{\psi}(z_{t},\,t,\,V(Y))$
8:  $z_{t+1}^{\star}\coloneqq H_{\theta}(z_{t},\,r_{t},\,V(Y),\,t)$// abduction
9:end for
[…]
13: Initialize $\bar{z}_{T}\coloneqq z_{T}^{\star}$
14:for $t=T$ down to 11 do
15:  $\bar{r}_{t}\coloneqq\ddot{\epsilon}_{\psi}(z_{t}^{\star},\,t,\,V(\bar{Y}))$
16:  $\bar{z}_{t-1}\coloneqq H_{\theta}^{-1}(\bar{z}_{t},\,\bar{r}_{t},\,V(\bar{Y}),\,t)$// prediction
17:end for
18:Return $\bar{x}=\textrm{Decode}(\bar{z}_{0})$
Appendix D Extended Ablation Results
@@ D.2 Effect of DDIM Steps
1. Ours⋆⋆$(T=50)$ 57.1 90.5 93.8 89.4 99.7 99.6 75.7 91.6 99.7 96.1 49.1 98 80.0 89.4 75.0 60
2. Ours⋆⋆$(T=100)$ 58.8 91.3 94.4 89.4 100 99.8 74.5 91.1 99.7 96.4 53.4 97.8 80.0 92.4 66.7 61.8
3. Ours⋆⋆$(T=200)$ 60.5 91.7 94.9 89.6 100 99.8 74.1 91.4 99.7 96.4 54.3 97.4 80.0 91.9 66.7 61.3
4. Ours⋆⋆$(T=500)$ 60.7 91.6 94.7 90.0 100 99.8 73.8 91.9 99.7 96.1 54.8 98.0 85.7 91.9 66.7 61.3
@@ D.3 Investigation of Attention Guidance
1. Ours 38.2 81.1 85.3 73.9 95.3 95.6 59.7 77.8 93.9 81.3 42.1 92.2 100 61.3 85.7 39.8
2. Ours 48.5 90.8 92.7 89.4 99 100 72.1 97 99.4 97.4 50.7 97.3 50 91.8 66.7 55.8
3. Ours⋆⋆ 58.8 91.3 94.4 89.4 100 99.8 74.5 91.1 99.7 96.4 53.4 97.8 80.0 92.4 66.7 61.8
@@
4. Ours 28.1 79.4 83.2 63.3 94.7 93.4 62.4 78.1 96.5 80.6 40.4 89.4 66.7 70.6 33.3 34.5
5. Ours 49.4 88.8 91.2 73.9 98.8 96.4 71.1 92.8 99.5 96.1 47.4 94.6 22.2 76.2 40 52.4
6. Ours⋆⋆ 57.1 88.7 94 77 99.4 100 73.3 90.4 99.7 95.6 49.4 95.1 66.7 80.7 50 54.4
@@
1. Ours 31.604 0.307
2. Ours 33.278 0.320
3. Ours⋆⋆ 31.52 0.312
@@
4. Ours 33.849 0.311
5. Ours⋆⋆ 31.885 0.299
6. Ours⋆⋆ 31.216 0.300
@@ D.4 Qualitative Evidence
Appendix E Additional Qualitative Results
@@ E.1 Pendulum
Figure 18: Pendulum counterfactuals from Causal-Adapter. $p$ for pendulum, $l$ for light, slsl for shadow length and spsp for shadow position.Traversal editing across four attributes demonstrates that our method produces high-quality, fine-grained generations of attribute variations.
Figure 19: Pendulum counterfactuals from Causal-Adapter. $p$ for pendulum, $l$ for light, slsl for shadow length and spsp for shadow position.

### v5

@@ Appendix A Extended Related Works [Back to Contents]
As shown in our motivational study (Appendix B), we evaluate the adaptation of text-to-image diffusion models for counterfactual image generation. Our findings show that relying solely on a frozen diffusion backbone with prompt-tuning is insufficient, as it fails to jointly represent causal semantic attributes and image embeddings. As a result, the model struggles to achieve precise counterfactual reasoning and generation. This highlights the need for an adaptive mechanism within the diffusion model, enhanced with injected causal semantics. Following the standard formulation of counterfactual image generation (Pearl, 2009; De Sousa Ribeiro et al., 2023; Wu et al., 2025; Rasal et al., 2025), where a predefined causal graph is treated as the structural prior, our Causal-Adapter models the SCM directly on the observed semantic attributes. This enforces correct causal dependencies during intervention and enables the generation of counterfactual images that are both visually plausible and causally faithful. When the SCM module is omitted, the framework naturally reduces to a standard conditional generation setting for that attribute, as illustrated in Figure 1.
@@
Appendix B Full Motivational Study Results [Back to Contents]
@@
3. Can existing T2I based methods do fine-grained editing? In causal editing, interventions on a single factor (e.g., increasing object size) should consistently induce predictable and proportional changes in the generated image. However, with current T2I based methods, numerical attributes must be mapped to discrete linguistic tokens, making it difficult to express edits along a continuous range. In practice, the primary way to adjust editing strength is by tuning the classifier-free guidance scale (Ho and Salimans, 2021). As shown in Figure 12, both NTI and MCPL fail to achieve fine-grained anatomical counterfactual editing of brain ventricular volume, even with different guidance scales. Their text-only control mechanisms cannot reliably translate numeric interventions into gradual visual changes. This limitation is particularly problematic in medical imaging domains, where precise, numerically controlled edits (e.g., adjusting brain or ventricular volume in MRI scans) are essential for simulating disease progression.
@@
Appendix C Dataset and Implementation [Back to Contents]
@@ C.1 Implementation Details
Unless specified, experimental setups on CelebA and ADNI follow the benchmark protocol in (Melistas et al., 2024), and CelebA-HQ experiments follow the setup in (Rasal et al., 2025). To ensure fairness, we adopt the same causal mechanism (normalizing flow) as described in the benchmark implementation. Experimental setup on Pendulum follows the configuration in CausalDiffAE (Komanduri et al., 2024c). Since both approaches assume linear causal modeling, we use our constructed simple MLP-based causal mechanism.
@@
For counterfactual image generation, we employ DDIM inversion (Song et al., 2021). To best preserve identity, we set the inversion guidance scale to $1.0$ (no classifier-free guidance (CFG)). For the forward process, we use CFG: CelebA uses $\alpha=3.0$ and CelebA-HQ uses $\alpha=1.5$ for effective edits, while ADNI and Pendulum perform well without CFG. Optionally, we incorporate token-level attention guidance manipulation following (Ju et al., 2024), replacing only the intervened token attentions while keeping the others fixed.
@@
CelebA ADNI Pendulum CelebA-HQ
Train Set Size 162,770 10,780 5,000 24000
Validation Set Size 19,867 0 500 3000
Test Set Size 19,962 2,240 2,000 3000
Resolution $\times 256\times 3 256×256$$\times 256\times 3$ $\times 256\times 3 256×256$$\times 256\times 3$
Downsampled Resolution $64\times 64\times 3$ $192\times 192\times 1$ $94\times 94\times 3$ $64\times 64\times 3$
Batch Size 40 40 40 40
Training Steps 200k 100k 50k 100k
Num of Attributes 4 6 4 7
Prompt Template ‘‘a human is @ and * and & and #’’ ‘‘an MRI image is @ and * and &’’ ‘‘an image is @ and * and & and #’’ ‘‘a human is @ and * and & and # and ! and ? and %’’
CTC Temperature $\tau$ 0.2 0.2 0.2 0.2
CTC Scale $\lambda$ 0.0005 0.0005 0.0005 0.0005
DDIM Steps $T$ 100 100 100 100
Guidance Scale $\alpha$ 3.0 1.0 1.0 1.5
@@ C.2 Metrics
Minimality evaluates whether a counterfactual differs from the factual image only in the intervened parent attribute, ideally leaving all other attributes unaffected. Counterfactual Latent Divergence (CLD) quantifies this by measuring the “distance” between counterfactual and factual images in a latent space (Sanchez and Tsaftaris, 2022). Intuitively, CLD captures a trade-off: the counterfactual should move sufficiently away from the factual class, but not farther than real samples belonging to the counterfactual class. Following Melistas et al. (2024), we compute CLD using an unconditional VAE. Specifically, we measure the KL divergence between the latent distributions of real and counterfactual images. The metric is minimized when both probabilities remain low, reflecting the balance between departing from the factual class while remaining closer to the counterfactual class than unrelated real samples.
@@ C.3 Full Regularization Algorithm
In the following, we present the fully regularized Causal-Adapter in Algorithms 12. Below we summarize the training and inference algorithm in high-level description.
Training:
  1. 1.
    Learn the causal mechanisms $F$: Using the predefined adjacency matrix $A$ derived from the causal graph and semantic attributes $Y$, the model learns how attributes causally influence one another. This produces a standalone SCM that governs how attributes should update when interventions are applied.
  2. 2.
    Train the conditional adapter: Each semantic attribute $y_{i}$ is projected into a token embedding $v_{i}(y_{i})$ through Prompt-Aligned Injection (PAI). With the diffusion model frozen, the adapter $\ddot{\epsilon}_{\psi}$ is trained to produce residuals $r_{t}$ conditioned on the attribute embeddings $V(Y)$. A diffusion loss $\mathcal{L}_{\text{DM}}$ ensures correct denoising behaviour, while a contrastive loss $\mathcal{L}_{\text{CTC}}$ encourages disentanglement and suppresses spurious attribute correlations.
Inference (Counterfactual Reasoning):
  1. 1.
    Abduction: The input image $x$ is inverted through the diffusion model (via DDIM inversion $H_{\theta}$) to obtain a latent trajectory $z_{0}^{\star},\ldots,z_{T}^{\star}]$ consistent with the observed image.
  2. 2.
    Action: A user-specified intervention $y_{i}^{\prime}$ is applied to the semantic attributes. The learned causal mechanisms update all causally connected attributes to produce a new attribute set $\bar{y}_{i}^{\prime}$ that respects the causal graph.
  3. 3.
    Prediction: Starting from the abducted latent $\bar{z}_{T}$, the model synthesizes a counterfactual image using the conditional adapter $\ddot{\epsilon}_{\psi}$ and the updated token embeddings $V\!\left(Y_{\mathrm{do}(y_{i}=y_{i}^{\prime})}\right)$, yielding the final counterfactual image $\bar{x}$.
@@
1:  Input: Image $x$, semantic attributes $Y=\{y_{i}\}_{i=1}^{K}$, binary adjacency matrix $A\in\{0,1\}^{K\times K}$, frozen modules $\{\mathcal{E},c_{\phi},\epsilon_{\theta}\}$, projector $G=\{g_{i}\}_{i=1}^{K}$
2:  Output: causal mechanisms $F=\{f_{i}\}_{i=1}^{K}$, updated placeholder embeddings $C=\{c_{i}\}_{i=1}^{K}$, updated projector $G=\{g_{i}\}_{i=1}^{K}$, causal adapter $\ddot{\epsilon}_{\psi}$.
3:  # Train causal mechanisms $F$
4:  $\bar{y}_{i}\coloneqq f_{i}(A_{i}\odot Y;\omega_{i})+u_{i},\quad u_{i}\sim\mathcal{N}(0,\sigma_{i}^{2}),$
5:  $F\coloneqq\arg\min_{F}\mathcal{L}_{\text{NLL}}(y_{i},\bar{y_{i}})$
6:  # Construct attribute injected token embeddings
7:  initialize placeholder embeddings $C=\{c_{i}\}_{i=1}^{K}$,
8:  $v_{i}(y_{i})=c_{i}+g_{i}(y_{i})$, $i=\{1,\cdots,K\}$
9:  $V(Y)=[v_{1}(y_{1}),\ldots,v_{K}(y_{K})]^{\top}$
10:  # Train conditional adapter
11:  for $t=1$ to $T$ do
12:  Encode latent: $z_{t}=\mathcal{E}(x,t)$
13:  Compute residual from causal adapter: $r_{t}=\ddot{\epsilon}_{\psi}(z_{t},t,V(Y))$
14:  Update parameters: $\psi,G,C\coloneqq\arg\min_{\psi,G,C}\left(\mathcal{L}_{\text{DM}}+\mathcal{L}_{\text{CTC}}\right)$
15:  end for
16:  Return $F=\{f_{i}\}_{i=1}^{K}$, $\ddot{\epsilon}_{\psi}$, $V$
Algorithm 2 Causal-Adapter: Inference
1:  Input: Image $x$, semantic attributes $Y$, frozen modules $\{\mathcal{E},c_{\phi},\epsilon_{\theta}\}$, trained causal mechanisms $F=\{f_{i}\}_{i=1}^{K}$, learned placeholder embedding $C=\{c_{i}\}_{i=1}^{K}$, trained causal adapter $\ddot{\epsilon_{\psi}}$, DDIM Inversion operator $H_{\theta}$ and its generative inverse $H_{\theta}^{-1}$
2:  Output: Counterfactual image $\bar{x}$
3:  $v_{i}(y_{i})=c_{i}+g_{i}(y_{i})$,   $i=\{1,\cdots,K\}$
4:  $V(Y)=[v_{1}(y_{1}),\ldots,v_{K}(y_{K})]^{\top}$
5:  # Abduction – infer inversed latent noise
6:  $z_{0}^{\star}=\mathcal{E}(x,0)$
7:  for $t=0$ to T1T-1 do
8:  $r_{t}=\ddot{\epsilon}_{\psi}(z_{t}^{\star},t,V(Y))$
9:  $z_{t+1}^{\star}=H_{\theta}(z_{t}^{\star},r_{t},V(Y),t)$
10:  end for
11:  # Action – apply intervention with $do(y_{i}=y_{i}^{\prime})$
12:  $\bar{y}_{i}^{\prime}\coloneqq f_{i}(A_{i}\odot Y_{do(y_{i}=y_{i}^{\prime})};\omega_{i})+u_{i},\quad u_{i}\sim\mathcal{N}(0,\sigma_{i}^{2}),\quad i=\{1,\cdots,K\}$
13:  $v_{i}(\bar{y}_{i}^{\prime})=c_{i}+g_{i}(\bar{y}_{i}^{\prime}),\quad i=\{1,\cdots,K\}$
14:  $V(Y_{do(y_{i}=y_{i}^{\prime})})=[v_{1}(\bar{y}_{1}^{\prime}),\ldots,v_{K}(\bar{y}_{K}^{\prime})]^{\top}$
15:  # Prediction – generate counterfactual
16:  initialize $\bar{z}_{T}\leftarrow z_{T}^{\star}$
17:  for $t=T$ to 11 do
18:  $\bar{r}_{t}=\ddot{\epsilon}_{\psi}(z_{t}^{\star},t,V(Y_{do(y_{i}=y_{i}^{\prime})}))$
19:  $\bar{z}_{t-1}=H_{\theta}^{-1}(\bar{z}_{t},\bar{r}_{t},V(Y_{do(y_{i}=y_{i}^{\prime})}),t)$
@@
21:  Return $\bar{x}$
Appendix D Extended Ablation Results [Back to Contents]
@@ D.2 Effect of DDIM Steps
2 Ours⋆⋆$(T=50)$ 57.1 90.5 93.8 89.4 99.7 99.6 75.7 91.6 99.7 96.1 49.1 98 80.0 89.4 75.0 60
3 Ours⋆⋆$(T=100)$ 58.8 91.3 94.4 89.4 100 99.8 74.5 91.1 99.7 96.4 53.4 97.8 80.0 92.4 66.7 61.8
4 Ours⋆⋆$(T=200)$ 60.5 91.7 94.9 89.6 100 99.8 74.1 91.4 99.7 96.4 54.3 97.4 80.0 91.9 66.7 61.3
5 Ours⋆⋆$(T=500)$ 60.7 91.6 94.7 90.0 100 99.8 73.8 91.9 99.7 96.1 54.8 98.0 85.7 91.9 66.7 61.3
@@ D.3 Investigation of Attention Guidance
3 Ours 38.2 81.1 85.3 73.9 95.3 95.6 59.7 77.8 93.9 81.3 42.1 92.2 100 61.3 85.7 39.8
4 Ours 48.5 90.8 92.7 89.4 99 100 72.1 97 99.4 97.4 50.7 97.3 50 91.8 66.7 55.8
5 Ours⋆⋆ 58.8 91.3 94.4 89.4 100 99.8 74.5 91.1 99.7 96.4 53.4 97.8 80.0 92.4 66.7 61.8
@@
7 Ours 28.1 79.4 83.2 63.3 94.7 93.4 62.4 78.1 96.5 80.6 40.4 89.4 66.7 70.6 33.3 34.5
8 Ours 49.4 88.8 91.2 73.9 98.8 96.4 71.1 92.8 99.5 96.1 47.4 94.6 22.2 76.2 40 52.4
9 Ours⋆⋆ 57.1 88.7 94 77 99.4 100 73.3 90.4 99.7 95.6 49.4 95.1 66.7 80.7 50 54.4
@@
3 Ours 31.604 0.307
4 Ours 33.278 0.320
5 Ours⋆⋆ 31.52 0.312
@@
7 Ours 33.849 0.311
8 Ours⋆⋆ 31.885 0.299
9 Ours⋆⋆ 31.216 0.300
@@ D.4 Qualitative Evidence
Appendix E Additional Results [Back to Contents]
@@ E.1 Pendulum
Figure 18: Pendulum counterfactuals from Causal-Adapter. $p$ for pendulum, $l$ for light, slsl for shadow length and $sp$ for shadow position.Traversal editing across four attributes demonstrates that our method produces high-quality, fine-grained generations of attribute variations.
Figure 19: Pendulum counterfactuals from Causal-Adapter. $p$ for pendulum, $l$ for light, slsl for shadow length and $sp$ for shadow position.
@@ E.2 CelebA
We report the full CelebA results with the corresponding means and standard deviations in Tables 1012. Compared with VAE, HVAE, and GAN, our method achieves best performance across most interventions, including up to an $86\%$ reduction in LPIPS (composition, from 0.122 to 0.017) and an $79\%$ reduction in FID (realism, from 27.861 to 8.152). Besides, our method delivers significant improvements under $do(a)$ and $do(br)$ on F1 score of bald attribute, demonstrating causal faithfulness in representing baldness across diverse individuals. As illustrated in Figure 20, Causal-Adapter can successfully add baldness to both male and female under $do(bl)$, and under interventions on causal parents such as $do(a)$, it can jointly edit age and baldness, whereas baselines fail to maintain causal consistency.
Table 12 additionally reports the LPIPS distance between each counterfactual and its original image, measuring identity preservation under different attribute interventions.
Table 10: Intervention effectiveness on CelebA test set. Age and Gender F1 under four interventions.
Method Age $(a)$ F1 $\uparrow$ Gender $(g)$ F1 $\uparrow$
$do(a)$ $do(g)$ $do(br)$ $do(bl)$ $do(a)$ $do(g)$ $do(br)$ $do(bl)$
2 VAE 35.00.04 78.20.02 81.60.02 81.90.02 97.70.01 90.90.02 95.90.02 97.30.01
3 HVAE 65.40.10 89.30.04 90.80.03 89.90.03 98.80.02 94.90.03 99.40.01 95.00.03
4 GAN 41.30.04 71.00.02 81.80.02 79.90.01 95.20.01 98.20.01 92.00.01 96.10.01
5 Ours 58.50.14 89.90.35 94.00.07 89.40.00 99.60.00 99.90.00 74.50.07 92.50.21
with attention guidance for localized editing
7 Ours (AG) 57.10.14 89.51.76 94.50.00 81.50.63 99.60.00 99.70.00 73.80.56 89.20.14
Table 11: Intervention effectiveness on CelebA test set. Beard and Bald F1 under four interventions.
Method Beard $(br)$ F1 $\uparrow$ Bald $(bl)$ F1 $\uparrow$
$do(a)$ $do(g)$ $do(br)$ $do(bl)$ $do(a)$ $do(g)$ $do(br)$ $do(bl)$
2 VAE 94.40.01 82.80.03 29.60.05 94.50.02 2.30.03 49.60.05 4.50.04 41.20.03
3 HVAE 95.20.03 95.10.03 44.10.11 91.60.04 2.00.05 86.00.05 4.50.07 61.10.04
4 GAN 90.80.01 83.80.02 23.30.03 90.70.01 2.10.02 82.00.02 5.50.02 49.20.02
5 Ours 99.70.00 96.10.00 52.10.21 98.10.07 59.20.91 91.20.28 74.52.75 58.80.35
with attention guidance for localized editing
7 Ours (AG) 99.70.07 96.80.00 48.20.14 96.60.07 38.71.90 78.40.49 47.71.76 51.40.49
[…]
Method Identity Preservation (LPIPS $\downarrow$) Composition Realism Minimality
Age Gender Bearded Bald MAE $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ CLD $\downarrow$
2 VAE 18.695 0.282 59.393 0.299
3 HVAE 7.143 0.122 35.712 0.305
4 GAN 60.120 0.276 27.861 0.304
5 Ours 0.0870.0066 0.1570.0067 0.07970.0060 0.1570.0073 3.5350.000 0.0170.000 8.1520.365 0.3100.001
Ours with attention guidance for improved realism and minimality
7 Ours (AG) 0.0610.0171 0.0720.0178 0.03300.0088 0.1090.0192 5.2130.173 0.3010.001